# Modeling Neural Activity with Conditionally Linear Dynamical Systems

### v1

@@ 4 Related work
Switching LDS models can be contrasted with models that smoothly interpolate between dynamical parameter regimes. The simplest example of this would be time-varying linear models (e.g. Costa et al. 2019); CLDS models are a generalization of this idea that comes with several advantages (see §2.2). Work by Costacurta et al. (2022) introduced an autoregressive model with a dynamics matrix that is subject to an approximately continuous and latent time warping factor. Unlike CLDS, this model does not infer a low-dimensional latent dynamical space. More similar to CLDS models is recent work by Hu et al. (2024). They relax the discrete switching in rSLDS models to allow smoothly varying soft mixtures of linear dynamics. Again, CLDS models achieve a similar effect but utilize observed experimental covariates to infer these dynamical transitions. Thus, CLDS models can loosely be seen as supervised analogs to these models.
@@ 5 Conclusion
In this work, we revisited and extended classical linear-Gaussian state space models of neural circuit dynamics. Our results suggest that these models can be competitive with modern methods when the dynamical parameters vary smoothly as a function of experimentally measured covariates. Like classical linear models, CLDS models are easy to fit and interpret. Our main technical contribution was to introduce an approximate GP prior over system parameters and show that this leads to closed form inference and parameter updates under a Gaussian noise model.
@@
As their name implies, CLDS models assume conditionally linear latent dynamics, and this assumption brings some limitations. First, these models rely on observing a time series of experimental covariates, ${\bm{u}}_{1:T}^{k}$. We expect performance to suffer if the covariates are corrupted for portions of time, such as during forecasting or with partial observations. Second, the model assumes linear dynamics conditioned on ${\bm{u}}_{t}^{k}$. We believe this is a good approximation in many settings of interest, particularly when there is strong tuning to sensory or behavioral variables—i.e. when the value of ${\bm{u}}_{t}^{k}$ can be used to accurately predict the position of ${\bm{x}}_{t}^{k}$. We expect (see §A.2) CLDS models to struggle in other settings where external measurements are only loosely correlated with the position of ${\mathbf{x}}_{t}^{k}$ (e.g. cognitive tasks with long periods of internal deliberation). In these scenarios, we expect that modern approaches that leverage deep learning (e.g. LFADS) will outperform CLDS models when given access to large amounts of data. Nevertheless, neural recordings are often trial-limited in practice (Williams & Linderman, 2021). We therefore view CLDS models as a broadly applicable modeling tool for many neuroscience applications.
@@
Future work could extend CLDS models to overcome these limitations, such as handling partially observed covariates, ${\bm{u}}_{t}^{k}$. Since CLDS models can be viewed as a dynamical extension of Wishart process models (see §4), future work could also apply this method to infer across-time noise correlations (reviewed in Panzeri et al., 2022), in addition to classical across-trial noise correlations. Nejatbakhsh et al. (2024) show how across-time correlations can be used to quantify similarity in dynamical systems—a topic that has recently attracted strong interest (Ostrow et al., 2023). CLDS models are a potentially attractive framework for tackling the unsolved challenge of estimating this high-dimensional correlation structure in trial-limited regimes.
@@ Acknowledgments
VG was supported by doctoral scholarships from the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Fonds de recherche du Québec – Nature et technologies (FRQNT). JWP was supported by grants from the Simons Collaboration on the Global Brain (SCGB AWD543027), the NIH BRAIN initiative (9R01DA056404-04), an NIH R01 (NIH 1R01EY033064), and a U19 NIH-NINDS BRAIN Initiative Award (U19NS104648). AHW was supported by the NIH BRAIN initiative (1RF1MH133778).
Impact Statement
The modeling presented here aims to provide a methodology to enhance our understanding of neural computation. Analysis of electrophysiological data can have long-term implications for the treatment and understanding of medical treatment and neurological disorders across different species. However, these considerations are far removed for the preliminary analyses and theoretical modeling presented, and we foresee no immediate societal consequences of this work.
@@ Appendix A Modeling
A.1 Expectation-Maximization Steps
A.1.1 Least squares derivation
@@
A.1.2 Joint dynamics and bias EM update in weight-space
@@
The quantity of interest for the M-step from the complete data log-likelihood is
@@
$=-\frac{1}{2}\mathbb{E}\left[\sum_{t=1}^{T-1}\left({\mathbf{x}}_{t+1}-\left({\mathbf{A}}({\bm{u}}_{t}){\mathbf{x}}_{t}+{\mathbf{b}}({\bm{u}}_{t})\right)\right)^{\top}Q^{-1}\left({\mathbf{x}}_{t+1}-\left({\mathbf{A}}({\bm{u}}_{t}){\mathbf{x}}_{t}+{\mathbf{b}}({\bm{u}}_{t})\right)\right)\right]$
@@
$\qquad\qquad+{\mathbf{x}}_{t}^{\top}{\mathbf{A}}({\bm{u}}_{t})^{\top}Q^{-1}{\mathbf{A}}({\bm{u}}_{t}){\mathbf{x}}_{t}+2{\mathbf{x}}_{t}^{\top}{\mathbf{A}}({\bm{u}}_{t})^{\top}Q^{-1}{\mathbf{b}}({\bm{u}}_{t})+{\mathbf{b}}({\bm{u}}_{t})^{\top}Q^{-1}{\mathbf{b}}({\bm{u}}_{t})\Bigg{]}$
@@
$\qquad\qquad+{\mathbf{A}}({\bm{u}}_{t})^{\top}Q^{-1}{\mathbf{A}}({\bm{u}}_{t})\mathbb{E}[{\mathbf{x}}_{t}{\mathbf{x}}_{t}^{\top}]+2{\mathbf{A}}({\bm{u}}_{t})^{\top}Q^{-1}{\mathbf{b}}({\bm{u}}_{t})\mathbb{E}[{\mathbf{x}}_{t}^{\top}]+{\mathbf{b}}({\bm{u}}_{t})^{\top}Q^{-1}{\mathbf{b}}({\bm{u}}_{t})\Bigg{]}$
Which with the function-space weights reads as
@@
$\frac{\partial\mathcal{L}}{\partial{\mathbf{W}}_{b}}$ $=\sum_{t=1}^{T-1}\bm{\phi}_{t}\mathbb{E}[{\mathbf{x}}_{t+1}^{\top}]Q^{-1}-(\bm{\phi}_{t}\bm{\phi}_{t}^{\top}\otimes\mathbb{E}[{\mathbf{x}}_{t}^{\top}]){\mathbf{W}}_{A}Q^{-1}+\bm{\phi}_{t}\bm{\phi}_{t}^{\top}{\mathbf{W}}_{b}Q^{-1}-{\mathbf{W}}_{b}$
@@
Incorporating the Gaussian prior on ${\mathbf{W}}_{A}$ and ${\mathbf{W}}_{b}$ and equating the two partial derivatives in (18-19) to 00 to obtain the stationary points, we obtain the system of equations
@@ A.2.1 CLDS conditional approximation error
$\mathbb{E}\left[|\mathcal{E}_{i}|\mid{\mathbf{u}}_{t}\right]\leq L_{i}\cdot\mathbb{E}\left[\left\|{\mathbf{x}}_{t}-{\bm{a}}\right\|_{2}^{2}\,\middle|\,{\mathbf{u}}_{t}\right].$ (26)
@@
$\mathbb{E}\left[|\mathcal{E}_{i}|\mid{\mathbf{u}}_{t}\right]\leq L_{i}\cdot\mathrm{Tr}\left[\mathbb{C}\mathrm{ov}[{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}]\right]$ (27)
@@
${\bm{A}}({\mathbf{u}}_{t})\coloneqq\nabla_{{\bm{x}}}\bm{f}(\mathbb{E}[{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}],{\mathbf{u}}_{t}),\quad{\bm{b}}({\mathbf{u}}_{t})\coloneqq\bm{f}(\mathbb{E}[{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}],{\mathbf{u}}_{t})-\nabla_{{\bm{x}}}\bm{f}(\mathbb{E}[{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}],{\mathbf{u}}_{t})\mathbb{E}[{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}].$ (30)
@@ A.2.2 Composite dynamics
$\bm{f}({\mathbf{x}}_{t},{\mathbf{u}}_{t})\approx\bm{g}({\mathbf{x}}_{t}).$ (31)
@@
$\mathbb{E}\left[\|\bm{f}({\mathbf{x}}_{t},{\mathbf{u}}_{t})-\bm{g}({\mathbf{x}}_{t})\|_{2}^{2}~{}\middle|~{}{\mathbf{x}}_{t}\right]$ (32)
@@
$\bm{g}({\mathbf{x}}_{t})=\mathbb{E}[\bm{f}({\mathbf{x}}_{t},{\mathbf{u}}_{t})\mid{\mathbf{x}}_{t}]=\mathbb{E}[{\mathbf{x}}_{t+1}\mid{\mathbf{x}}_{t}]$ (33)
@@
$\mathcal{R}_{i}$ $=f_{i}({\mathbf{x}}_{t},{\mathbf{u}}_{t})-\mathbb{E}[h_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})\mid{\mathbf{x}}_{t}]$ (35)
@@
$\mathcal{R}_{i}$ $=\mathbb{E}[f_{i}({\mathbf{x}}_{t},{\mathbf{u}}_{t})-h_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})\mid{\mathbf{x}}_{t}]$ (36)
$=\mathbb{E}[f_{i}({\mathbf{x}}_{t},{\mathbf{u}}_{t})-f_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})+f_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})-h_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})\mid{\mathbf{x}}_{t}]$ (37)
@@
$|\mathcal{R}_{i}|$ $\leq\mathbb{E}_{\tilde{{\mathbf{u}}}_{t}\mid{\mathbf{x}}_{t}}[\,\big{|}f_{i}({\mathbf{x}}_{t},{\mathbf{u}}_{t})-f_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})+f_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})-h_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})\big{|}\,]$ (38)
$\leq\mathbb{E}_{\tilde{{\mathbf{u}}}_{t}\mid{\mathbf{x}}_{t}}[\,\big{|}f_{i}({\mathbf{x}}_{t},{\mathbf{u}}_{t})-f_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})\big{|}\,]+\mathbb{E}_{\tilde{{\mathbf{u}}}_{t}\mid{\mathbf{x}}_{t}}[\,\big{|}f_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})-h_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})\big{|}\,]$ (39)
@@
$|f_{i}({\mathbf{x}}_{t},{\mathbf{u}})-f_{i}({\mathbf{x}}_{t},{\mathbf{u}}^{\prime})|\leq C_{i}\|{\mathbf{u}}-{\mathbf{u}}^{\prime}\|_{2}\qquad\forall{\mathbf{u}},{\mathbf{u}}^{\prime}\in\mathcal{U}.$ (41)
@@
$\mathbb{E}_{{\mathbf{u}}_{t},\tilde{{\mathbf{u}}}_{t}\mid{\mathbf{x}}_{t}}[\,\big{|}f_{i}({\mathbf{x}}_{t},{\mathbf{u}}_{t})-f_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})\big{|}\,]$ $\leq C_{i}\cdot\mathbb{E}_{{\mathbf{u}}_{t},\tilde{{\mathbf{u}}}_{t}\mid{\mathbf{x}}_{t}}[\,\big{\|}{\mathbf{u}}_{t}-\tilde{{\mathbf{u}}}_{t}\big{\|}_{2}\,]$ (42)
$\leq C_{i}\sqrt{\mathbb{E}_{{\mathbf{u}}_{t},\tilde{{\mathbf{u}}}_{t}\mid{\mathbf{x}}_{t}}\big{\|}{\mathbf{u}}_{t}-\tilde{{\mathbf{u}}}_{t}\big{\|}_{2}^{2}}$ (43)
$=C_{i}\sqrt{2\operatorname{Tr}[\mathbb{C}\mathrm{ov}[{\mathbf{u}}_{t}\mid{\mathbf{x}}_{t}]]}$ (44)
@@
$\mathbb{E}_{{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}}[\,\big{|}f_{i}({\mathbf{x}}_{t},{\mathbf{u}}_{t})-h_{i}({\mathbf{x}}_{t},{\mathbf{u}}_{t})\big{|}\,]\leq L_{i}\cdot\operatorname{Tr}[\mathbb{C}\textrm{ov}[{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}]]$ (45)
@@
$\mathbb{E}|\mathcal{R}_{i}|\leq C_{i}\cdot\mathbb{E}_{{\mathbf{x}}_{t}}\sqrt{2\operatorname{Tr}[\mathbb{C}\mathrm{ov}[{\mathbf{u}}_{t}\mid{\mathbf{x}}_{t}]]}+L_{i}\cdot\mathbb{E}_{{\mathbf{u}}_{t}}\operatorname{Tr}[\mathbb{C}\textrm{ov}[{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}]]$ (46)
@@
$\sum_{i=1}^{D}\mathbb{E}|\mathcal{R}_{i}|\leq C\cdot\mathbb{E}_{{\mathbf{x}}_{t}}\sqrt{2\operatorname{Tr}[\mathbb{C}\mathrm{ov}[{\mathbf{u}}_{t}\mid{\mathbf{x}}_{t}]]}+L\cdot\mathbb{E}_{{\mathbf{u}}_{t}}\operatorname{Tr}[\mathbb{C}\textrm{ov}[{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}]]$ (47)

### v2

@@ 4 Related work
Switching LDS models can be contrasted with models that smoothly interpolate between dynamical parameter regimes. The simplest example of this would be time-varying linear models (e.g. costa2019adaptive ); CLDS models are a generalization of this idea that comes with several advantages (see §2.2). Recent work (costacurta2022distinguishing, ) relaxed switching dynamics to be subject to an approximately continuous and latent time warping factor. More similar to CLDS models are smoothly varying latent linear dynamical models, such as the gpSLDS (hu2024modeling, ), which relax the discrete switching in rSLDS models to allow smoothly varying soft mixtures of linear dynamics, and the dLDS (JMLR:v25:23-0777, ), which learns a dictionary of linear dynamics combined with time-varying coefficients. Again, the primary difference is that CLDS models achieve a similar effect but utilize observed experimental covariates to infer these dynamical transitions. This approach is closer to previous work (pmlr-v70-foerster17a, ) using neural network architectures to parametrize input-dependent linear recurrent dynamics.
@@ 5 Conclusion
In this work, we extend classical linear-Gaussian state space models of neural dynamics. Our results suggest that these models are competitive with modern methods when the dynamical parameters vary smoothly as a function of available covariates. Like classical linear models, CLDS models are easy to fit and interpret. Our main technical contribution was to introduce an approximate GP prior over system parameters and show that this leads to closed-form inference under a Gaussian noise model.
While these results are promising, CLDS models have limitations that merit consideration. To start, we did not extend to non-Gaussian observations to model spike counts directly, nor leveraged the prior structure on the coefficients $\mathbf{F}$} to convey more complete parameter posteriors—we discuss both of these options in the text, and see them as essential avenues for future work. Furthermore, as their name implies, CLDS models assume conditionally linear latent dynamics. We detail in Appendix §A.2 how our approach can be viewed as a first-order (in $\mathbf{x}$}) approximation to a full nonlinear system of the form
$\mathbb{E}[{\mathbf{x}}_{t+1}\mid{\mathbf{x}}_{t},{\bm{u}}_{t}]=\bm{f}_{x}({\mathbf{x}}_{t},{\bm{u}}_{t}),\quad\mathbb{E}[{\mathbf{y}}_{t}\mid{\mathbf{x}}_{t},{\bm{u}}_{t}]=\bm{f}_{y}({\mathbf{x}}_{t},{\bm{u}}_{t}).$
@@
We derive upper bounds for the approximation error between our linearized dynamics and this nonlinear system, focusing of $\bm{f}_{x}$ and showing that the quality of this approximation is guaranteed to be small if $\bm{f}_{x}$ is well-behaved and the conditional covariance in ${\mathbf{x}}_{t}$ given ${\bm{u}}_{t}$ is small. These same guarantees help shed light onto limitations. First, these models rely on observing a time series of experimental covariates ${\bm{u}}_{1:T}$ and performance would suffer if the covariates are corrupted, such as during forecasting or with partial observations. Second, the model assumes linear dynamics conditioned on ${\bm{u}}_{t}$. We believe this is a good approximation in many settings of interest, particularly when there is strong tuning in $\mathbf{x}$} to sensory or behavioral variables $\mathbf{u}$}, and expect CLDS models to struggle when they are loosely correlated (e.g. cognitive tasks with long periods of internal deliberation). In these scenarios, we expect that more nonlinear approaches will outperform CLDS models when given access to large amounts of data. Nevertheless, neural recordings are often trial-limited in practice (williams2021statistical, ). We therefore view CLDS models as a broadly applicable modeling tool for many neuroscience applications.
@@
Future work could extend CLDS models to overcome these limitations, such as handling partially observed covariates, ${\bm{u}}_{t}^{k}$. Since CLDS models can be viewed as a dynamical extension of Wishart process models (see §4), future work could also apply this method to infer across-time noise correlations (reviewed in panzeri2022structures, ), in addition to classical across-trial noise correlations. Recent work (nejatbakhsh2025comparing, ) shows how across-time correlations can be used to quantify similarity in dynamical systems—a topic that has recently attracted strong interest (ostrow2023beyond, ). CLDS models are a potentially attractive framework for tackling the unsolved challenge of estimating this high-dimensional correlation structure in trial-limited regimes.
@@ Acknowledgments
VG was supported by the Porter Ogden Jacobus Fellowship at Princeton University and by doctoral scholarships from the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Fonds de recherche du Québec – Nature et technologies (FRQNT). JWP was supported by grants from the Simons Collaboration on the Global Brain (SCGB AWD543027), the NIH BRAIN initiative (9R01DA056404-04), an NIH R01 (NIH 1R01EY033064), and a U19 NIH-NINDS BRAIN Initiative Award (U19NS104648). AHW was supported by the NIH BRAIN initiative (1RF1MH133778).
@@ Appendix A Modeling
A.1 Inference
A.1.1 Basis space form
We want to show
${\mathbf{M}}({\bm{u}}){\bm{X}}={\mathbf{W}}^{\top}(\bm{\phi}({\bm{u}})\otimes{\bm{X}})$
for $\mathbf{M}$} defined as in (1c) with basis functions $\{\phi_{\ell}\}_{\ell=1}^{L}$, ${\bm{u}}\in\mathcal{U}$, and ${\bm{X}}\in\mathbb{R}^{D_{2}\times D_{3}}$, $D_{3}\in\mathbb{N}$.
For each $i,j$-th entry,
$\left[{\mathbf{M}}({\bm{u}}){\bm{X}}\right]_{ij}$ $=\sum_{k=1}^{D_{2}}\left[{\mathbf{M}}({\bm{u}})\right]_{ik}\left[{\bm{X}}\right]_{kj}$
$=\sum_{k=1}^{D_{2}}\sum_{\ell=1}^{L}{\textnormal{w}}^{(ik)}_{\ell}\phi_{\ell}({\bm{u}}){\bm{X}}_{kj}$
$=\sum_{k=1}^{D_{2}}\sum_{\ell=1}^{L}{\textnormal{w}}^{(ik)}_{\ell}\left(\bm{\phi}({\bm{u}})\otimes{\bm{X}}\right)_{D_{2}\ell+k,j}$
$=\sum_{k=1}^{D_{2}}\sum_{\ell=1}^{L}{\mathbf{W}}^{\top}_{i,D_{2}\ell+k}\left(\bm{\phi}({\bm{u}})\otimes{\bm{X}}\right)_{D_{2}\ell+k,j}$
$={\mathbf{W}}^{\top}_{i,:}\left(\bm{\phi}({\bm{u}})\otimes{\bm{X}}\right)_{:,j}$
$=\left[{\mathbf{W}}^{\top}\left(\bm{\phi}({\bm{u}})\otimes{\bm{X}}\right)\right]_{ij}$
as desired, where we’ve defined ${\mathbf{W}}_{D_{2}\ell+k,i}\coloneqq{\textnormal{w}}^{(ik)}_{\ell}$.
A.1.2 Least squares derivation for Conditionally Linear Regression
@@
A.1.3 Joint dynamics and bias EM update in weight-space
@@
The quantity of interest for the M-step from the complete data log-likelihood is (up to an additive constant)
@@
$=-\frac{1}{2}\mathbb{E}\left[\sum_{t=1}^{T-1}\left({\mathbf{x}}_{t+1}-\left({\mathbf{A}}({\bm{u}}_{t}){\mathbf{x}}_{t}+{\mathbf{b}}({\bm{u}}_{t})\right)\right)^{\top}Q^{-1}\left({\mathbf{x}}_{t+1}-\left({\mathbf{A}}({\bm{u}}_{t}){\mathbf{x}}_{t}+{\mathbf{b}}({\bm{u}}_{t})\right)\right)\right]+\textrm{const.}$
@@
$\qquad\qquad+{\mathbf{x}}_{t}^{\top}{\mathbf{A}}({\bm{u}}_{t})^{\top}Q^{-1}{\mathbf{A}}({\bm{u}}_{t}){\mathbf{x}}_{t}+2{\mathbf{x}}_{t}^{\top}{\mathbf{A}}({\bm{u}}_{t})^{\top}Q^{-1}{\mathbf{b}}({\bm{u}}_{t})+{\mathbf{b}}({\bm{u}}_{t})^{\top}Q^{-1}{\mathbf{b}}({\bm{u}}_{t})\Bigg]+\textrm{const.}$
@@
$\qquad\qquad+{\mathbf{A}}({\bm{u}}_{t})^{\top}Q^{-1}{\mathbf{A}}({\bm{u}}_{t})\mathbb{E}[{\mathbf{x}}_{t}{\mathbf{x}}_{t}^{\top}]+2{\mathbf{A}}({\bm{u}}_{t})^{\top}Q^{-1}{\mathbf{b}}({\bm{u}}_{t})\mathbb{E}[{\mathbf{x}}_{t}^{\top}]+{\mathbf{b}}({\bm{u}}_{t})^{\top}Q^{-1}{\mathbf{b}}({\bm{u}}_{t})\Bigg]+\textrm{const.}$
Which with the finite basis expansion reads as
@@
$\frac{\partial\mathcal{L}}{\partial{\mathbf{W}}_{b}}$ $=\sum_{t=1}^{T-1}\bm{\phi}_{t}\mathbb{E}[{\mathbf{x}}_{t+1}^{\top}]Q^{-1}-(\bm{\phi}_{t}\bm{\phi}_{t}^{\top}\otimes\mathbb{E}[{\mathbf{x}}_{t}^{\top}]){\mathbf{W}}_{A}Q^{-1}+\bm{\phi}_{t}\bm{\phi}_{t}^{\top}{\mathbf{W}}_{b}Q^{-1}$
@@
Incorporating the Gaussian prior on ${\mathbf{W}}_{A}$ and ${\mathbf{W}}_{b}$ and equating the two partial derivatives in (14-15) to 0 to obtain the stationary points, we obtain the system of equations
@@ A.2.1 CLDS conditional approximation error
$\mathbb{E}\left[|\mathcal{E}_{i}|\mid{\mathbf{u}}_{t}\right]\leq L_{i}\cdot\mathbb{E}\left[\left\|{\mathbf{x}}_{t}-{\bm{a}}\right\|_{2}^{2}\,\middle|\,{\mathbf{u}}_{t}\right].$
@@
$\mathbb{E}\left[|\mathcal{E}_{i}|\mid{\mathbf{u}}_{t}\right]\leq L_{i}\cdot\mathrm{Tr}\left[\mathbb{C}\mathrm{ov}[{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}]\right]$
@@
${\bm{A}}({\mathbf{u}}_{t})\coloneqq\nabla_{{\bm{x}}}\bm{f}(\mathbb{E}[{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}],{\mathbf{u}}_{t}),\quad{\bm{b}}({\mathbf{u}}_{t})\coloneqq\bm{f}(\mathbb{E}[{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}],{\mathbf{u}}_{t})-\nabla_{{\bm{x}}}\bm{f}(\mathbb{E}[{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}],{\mathbf{u}}_{t})\mathbb{E}[{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}].$
@@ A.2.2 Composite dynamics
$\bm{f}({\mathbf{x}}_{t},{\mathbf{u}}_{t})\approx\bm{g}({\mathbf{x}}_{t}).$
@@
$\mathbb{E}\left[\|\bm{f}({\mathbf{x}}_{t},{\mathbf{u}}_{t})-\bm{g}({\mathbf{x}}_{t})\|_{2}^{2}~\middle|~{\mathbf{x}}_{t}\right]$
@@
$\bm{g}({\mathbf{x}}_{t})=\mathbb{E}[\bm{f}({\mathbf{x}}_{t},{\mathbf{u}}_{t})\mid{\mathbf{x}}_{t}]=\mathbb{E}[{\mathbf{x}}_{t+1}\mid{\mathbf{x}}_{t}]$
@@
$\mathcal{R}_{i}$ $=f_{i}({\mathbf{x}}_{t},{\mathbf{u}}_{t})-\mathbb{E}[h_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})\mid{\mathbf{x}}_{t}]$
@@
$\mathcal{R}_{i}$ $=\mathbb{E}[f_{i}({\mathbf{x}}_{t},{\mathbf{u}}_{t})-h_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})\mid{\mathbf{x}}_{t}]$
$=\mathbb{E}[f_{i}({\mathbf{x}}_{t},{\mathbf{u}}_{t})-f_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})+f_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})-h_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})\mid{\mathbf{x}}_{t}]$
@@
$|\mathcal{R}_{i}|$ $\leq\mathbb{E}_{\tilde{{\mathbf{u}}}_{t}\mid{\mathbf{x}}_{t}}[\,\big|f_{i}({\mathbf{x}}_{t},{\mathbf{u}}_{t})-f_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})+f_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})-h_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})\big|\,]$
$\leq\mathbb{E}_{\tilde{{\mathbf{u}}}_{t}\mid{\mathbf{x}}_{t}}[\,\big|f_{i}({\mathbf{x}}_{t},{\mathbf{u}}_{t})-f_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})\big|\,]+\mathbb{E}_{\tilde{{\mathbf{u}}}_{t}\mid{\mathbf{x}}_{t}}[\,\big|f_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})-h_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})\big|\,]$
@@
$|f_{i}({\mathbf{x}}_{t},{\mathbf{u}})-f_{i}({\mathbf{x}}_{t},{\mathbf{u}}^{\prime})|\leq C_{i}\|{\mathbf{u}}-{\mathbf{u}}^{\prime}\|_{2}\qquad\forall{\mathbf{u}},{\mathbf{u}}^{\prime}\in\mathcal{U}.$
@@
$\mathbb{E}_{{\mathbf{u}}_{t},\tilde{{\mathbf{u}}}_{t}\mid{\mathbf{x}}_{t}}[\,\big|f_{i}({\mathbf{x}}_{t},{\mathbf{u}}_{t})-f_{i}({\mathbf{x}}_{t},\tilde{{\mathbf{u}}}_{t})\big|\,]$ $\leq C_{i}\cdot\mathbb{E}_{{\mathbf{u}}_{t},\tilde{{\mathbf{u}}}_{t}\mid{\mathbf{x}}_{t}}[\,\big\|{\mathbf{u}}_{t}-\tilde{{\mathbf{u}}}_{t}\big\|_{2}\,]$
$\leq C_{i}\sqrt{\mathbb{E}_{{\mathbf{u}}_{t},\tilde{{\mathbf{u}}}_{t}\mid{\mathbf{x}}_{t}}\big\|{\mathbf{u}}_{t}-\tilde{{\mathbf{u}}}_{t}\big\|_{2}^{2}}$
$=C_{i}\sqrt{2\operatorname{Tr}[\mathbb{C}\mathrm{ov}[{\mathbf{u}}_{t}\mid{\mathbf{x}}_{t}]]}$
@@
$\mathbb{E}_{{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}}[\,\big|f_{i}({\mathbf{x}}_{t},{\mathbf{u}}_{t})-h_{i}({\mathbf{x}}_{t},{\mathbf{u}}_{t})\big|\,]\leq L_{i}\cdot\operatorname{Tr}[\mathbb{C}\textrm{ov}[{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}]]$
@@
$\mathbb{E}|\mathcal{R}_{i}|\leq C_{i}\cdot\mathbb{E}_{{\mathbf{x}}_{t}}\sqrt{2\operatorname{Tr}[\mathbb{C}\mathrm{ov}[{\mathbf{u}}_{t}\mid{\mathbf{x}}_{t}]]}+L_{i}\cdot\mathbb{E}_{{\mathbf{u}}_{t}}\operatorname{Tr}[\mathbb{C}\textrm{ov}[{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}]]$
@@
$\sum_{i=1}^{D}\mathbb{E}|\mathcal{R}_{i}|\leq C\cdot\mathbb{E}_{{\mathbf{x}}_{t}}\sqrt{2\operatorname{Tr}[\mathbb{C}\mathrm{ov}[{\mathbf{u}}_{t}\mid{\mathbf{x}}_{t}]]}+L\cdot\mathbb{E}_{{\mathbf{u}}_{t}}\operatorname{Tr}[\mathbb{C}\textrm{ov}[{\mathbf{x}}_{t}\mid{\mathbf{u}}_{t}]]$